OpenStudy (writerforever18):

Scott invested his savings in two investment funds. The $8000 that he invested in Fund A returned a 7% profit. The amount that he invested in Fund B returned a 2% profit. How much did he invest in Fund B, if both funds together returned a 4% profit?

jimthompson5910 (jim_thompson5910):

`The $8000 that he invested in Fund A returned a 7% profit` this means that Scott earned 0.07*8000 = 560 dollars in profit from fund A alone ---------------------------------------- `The amount that he invested in Fund B returned a 2% profit` meaning that he earned 0.02x dollars from Fund B alone where x is the amount invested in Fund B. It's unknown for now. ---------------------------------------- In total, the profits from A and B combine to 560+0.02x dollars Is this making sense so far?

OpenStudy (writerforever18):

Yes, it makes perfect sense! :)

jimthompson5910 (jim_thompson5910):

ok great

jimthompson5910 (jim_thompson5910):

now we know $8000 was invested in fund A x dollars is put into fund B in total, he put 8000+x dollars into both funds ---------------------------------- we are told that `both funds together returned a 4% profit` meaning that you take 4% of the total and that's the profit made ie, 4% of total = 0.04*(8000+x) = total profit made

jimthompson5910 (jim_thompson5910):

earlier I showed how 560+0.02x dollars was also the total profit the two total profits are the same, so equate them 560+0.02x = 0.04*(8000+x) now solve for x to get your answer

OpenStudy (writerforever18):

Awesome thank you so much!!! :) Just curious, where did 560 come from?

jimthompson5910 (jim_thompson5910):

the profit made in fund A 7% of 8000 = 0.07*8000 = 560

OpenStudy (writerforever18):

Oh wait a minute...I just went back up and found that out haha. But again, thank you so much for your help!

jimthompson5910 (jim_thompson5910):

that's ok and you're welcome

OpenStudy (writerforever18):

Does the answer 12,000 sound right to you?

jimthompson5910 (jim_thompson5910):

12,000 dollars is the amount invested in fund B, correct

OpenStudy (writerforever18):

Yes

OpenStudy (writerforever18):

Ok it's ok it was right! Thanks! :)

OpenStudy (writerforever18):

Theres one more, would you be able to help me please?

jimthompson5910 (jim_thompson5910):

sure go ahead